OpenStudy (anonymous):

Find the equation of a line through (4, -3) that is parallel to the line that passes through these points (-4,7) and (2,4). Write your answer in standard form (Ax + By = C). Show all your work.

OpenStudy (zzr0ck3r):

if its parallel to the line that passes through (-4,7) and (2,4), then it has the same slope so find slope (4-7)/(2-(-4)) = -3/6 = -1/2 so we know the new line looks like \[y = (-1/2)x+b\] so we use the first given point (4, -3) to figure out b plug in 4 for x and -3 for y -3 = (-1/2)4+b ---> -3=-2+b so b = -1 so you have \[y=-\frac{1}{2}x-1\]

OpenStudy (zzr0ck3r):

understand?